The African nation Mauritius estimates that in 2015 they had 584,600 people in their labor force, of whom 538,300 were employed. That means 46,300 people were unemployed. Additionally, 383,700 adults were not in the labor force. Calculate the labor force participation rate to the nearest hundredth. (Hint: Answer is percentage terms, so a 50% LFPR rate would be 50.00, not 0.50. Calculate the LFPR as (E+U)/(E+U+NILF)*100)
Answer #1

Given Labor Force = 584,600

People who are not in the labor force (adult population) = 383,700

Therefore, total adult population = 584,600+383,700= 968,300

Labor Force Participation Rate = (Labour Force / Adult Population)*100

LFPR = (584,600/968,300)*100

LFPR = 60.37%

The labor force participation rate = 60.37 (ignoring percentage sign)